# Continual Learning on CLIP via Incremental Prompt Tuning with Intrinsic Textual Anchors

### v2

@@ Abstract
Continual learning (CL) enables deep neural networks to acquire new knowledge over time while mitigating catastrophic forgetting of previously learned information. The powerful generalization ability of pre-trained models (PTMs), such as the Contrastive Language-Image Pre-training (CLIP) model, has inspired a range of CL methods targeting new and specialized tasks, further bridging the gap between PTMs and continual adaptation. Leveraging its multi-modal visual and textual representations, CLIP offers a natural paradigm for CL, where new tasks can be accommodated by incrementally learning lightweight parameters, particularly prompts. However, existing prompt-based CL methods for PTMs often rely on complex designs built upon specific assumptions, such as intricate regularization schemes for prompt pools, specialized routing mechanisms, or multi-stage incrementation processes. While these approaches improve performance, they frequently introduce additional—and possibly unnecessary—complexity, underutilizing CLIP’s intrinsic capabilities. In this paper, we propose a concise CL approach for CLIP based on incremental prompt tuning that fully exploits its multi-modal structure and the stability of textual representations. Our method, Textual Prototype-guided Prompt Tuning (TPPT), introduces textual prototypes not merely as static classifiers, as in existing methods, but as stable anchors to guide the learning of visual prompts, thereby shaping the embedding space (i.e., TPPT-V). We show that our bidirectional supervision strategy enables more effective learning of new knowledge while reducing forgetting. To further close the vision-language gap during CL, we activate the language branch and extend our approach to jointly optimize both visual and textual prompts (i.e., TPPT-VT). We also introduce a relational diversity regularization on the textual anchors to prevent embedding space collapse and mitigate correlated forgetting. Extensive experiments and analyses demonstrate the effectiveness of our proposed approach, highlighting the benefits of leveraging CLIP’s intrinsic guidance for continual adaptation.
@@ 1 Introduction
Continual learning (CL) focuses on learning from a sequence of data or tasks while accumulating knowledge over time and avoiding catastrophic forgetting of previously learned information. It aims to do so with minimal storage and computational overhead by eliminating the need to retain all previously encountered data (De Lange et al., 2021; Gomes et al., 2017; Mai et al., 2022). A representative CL setting is class-incremental learning (CIL), where classification tasks arrive incrementally, each consisting of distinct, non-overlapping classes (Zhou et al., 2023a; Rebuffi et al., 2017). To mitigate catastrophic forgetting, various CL approaches have been explored, including experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022) and regularization-based techniques (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023), both of which have shown notable success.
@@
Recent developments in pre-trained models (PTMs) based on the Transformer architecture (Vaswani et al., 2017) have demonstrated strong generalization capabilities and achieved state-of-the-art performance across various applications (Dosovitskiy et al., 2020; Devlin et al., 2018; Wortsman et al., 2022). Among these, multi-modal pre-trained models, such as Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021), jointly learn vision-language embedding spaces, effectively aligning visual and textual representations and enabling robust zero-shot generalization. To better adapt PTMs to downstream tasks, parameter-efficient fine-tuning methods such as prompt tuning (Zhou et al., 2022c; Yao et al., 2023; Khattak et al., 2023a; b) and adapters (Gao et al., 2023; Zhang et al., 2021; Sung et al., 2022) have been developed.
In CL settings, PTMs such as Vision Transformer (ViT) (Dosovitskiy et al., 2020) and CLIP (Radford et al., 2021) often serve as stable, “frozen” backbones for continually arriving downstream tasks (e.g., in CIL). CL methods built on PTMs incrementally learn a fixed or dynamically expanding set of trainable parameters, such as soft prompts (Wang et al., 2022e; d; Lester et al., 2021; Liu et al., 2021; Smith et al., 2023) or adapters (Jha et al., 2024; Zhou et al., 2023b; McDonnell et al., 2024; Wang et al., 2024; Yu et al., 2024), to incorporate new tasks encountered in the data stream. CLIP, with its jointly trained vision and language representations, provides a natural foundation for CL (Jha et al., 2024; Wang et al., 2023), where textual embeddings can serve as pre-trained universal classifier heads for visual tasks. Due to their lightweight and non-intrusive nature, prompt-based methods have gained popularity in CL, particularly for CIL.
However, existing prompt-based continual learning methods with pre-trained models (including CLIP and ViT) often rely on task-specific prompt expansion (Smith et al., 2023; Wang et al., 2022d), complex regularization using external priors like orthogonality (Smith et al., 2023; Wang et al., 2022e; 2023), or intricate prompt routing schemes (Smith et al., 2023; Wang et al., 2022d; c). While these methods can boost performance, they often introduce additional and potentially unnecessary complexity, potentially underutilizing the intrinsic strengths of CLIP’s multi-modal representations and the stability of its textual embeddings.
@@
Figure 1: Conceptual illustrations of: (a) standard Cross-Entropy (CE), (b) our proposed TPPT-V, (c) a naïve multi-modal extension of TPPT-V, and (d) our proposed TPPT-VT. (a) Prior methods (Zhou et al., 2022c; Jia et al., 2022; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2023; Zhou et al., 2023c) use CE loss to adapt PTMs, but suffer from representation drift (Gama et al., 2014; Lu et al., 2018), leading to forgetting. (b) TPPT-V introduces a textual prototypical contrastive loss to anchor visual features and mitigate drift. (c) A naïve extension that also tunes textual prompts may improve textual prototype quality but risks collapse to trivial solutions (Minderer et al., 2022; Kim et al., 2022; Khattak et al., 2023a; Liang et al., 2022). (d) TPPT-VT addresses this by regularizing multi-modal prompt learning with diversity constraints on textual prototypes.
@@
It accommodates new tasks by incrementally adding visual prompts (Jia et al., 2022) to the inputs of multiple multi-head self-attention (MSA) layers within the CLIP encoders. While prompt tuning has been widely used for CL with PTMs such as CLIP (Jha et al., 2024) and ViT (Smith et al., 2023; Wang et al., 2022d), prior works often introduce complexity through orthogonal regularization (Smith et al., 2023), probabilistic modeling (Jha et al., 2024), or handcrafted prompt selection (Smith et al., 2023; Wang et al., 2023).
@@
We first present TPPT-V, which applies incremental prompt tuning only to the vision encoder of CLIP, guided by fixed textual embeddings of class labels. Prior works (Wang et al., 2023; Thengane et al., 2022; Ding et al., 2022; Zheng et al., 2023) typically treat these textual embeddings as static classifiers and use cross entropy (CE) loss to align vision features with corresponding textual embeddings. While this is effective when all classes are jointly trained, relying solely on CE leads to representation drift and interference (Gama et al., 2014; Lu et al., 2018) in CL due to partial class visibility and the model’s unawareness of future tasks—resulting in forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989), as illustrated in Fig. 1(a) and Fig. LABEL:fig:intro_drift.
As conceptualized in Fig. 1(a), this issue partly arises because CE loss focuses only on assigning each sample in isolation to a class label, ignoring how the embedding space evolves over time. We show that treating the fixed textual embeddings as prototype anchors (Li et al., 2020; Snell et al., 2017; Lu et al., 2024) and reintroducing contrastive learning supervision (Radford et al., 2021; Khosla et al., 2020; Grill et al., 2020) alleviates this problem. Although the textual prototypes remain fixed, the contrastive loss (specifically, the asymmetric CE formulation) serves as an additional regularizer, guiding visual embeddings away from unrelated prototypes and stabilizing the embedding space (Fig. 1(b)). In this way, CLIP’s stable textual prototypes serve as anchors that help preserve prior knowledge by regularizing representation drift in the embedding space during CL.
To further enhance vision-language alignment in CL on downstream tasks, we go beyond vision-only tuning and enable learning of both visual and textual prompts. However, naïve joint prompt tuning can cause embedding space collapse (Minderer et al., 2022; Kim et al., 2022; Liang et al., 2022; Khattak et al., 2023a) (as shown in Fig. 1(c) and Fig. LABEL:fig:intro_collapse), especially in CL, where the model can easily converge to trivial solutions (Minderer et al., 2022; Kim et al., 2022; Liang et al., 2022; Khattak et al., 2023a). To prevent this, we propose TPPT-VT (illustrated in Fig. 1(d)), which introduces relational diversity regularization on textual prompts, maintaining diversity among textual prototypes and mitigating collapse. By preserving the original contextual structure of textual representations, TPPT-VT improves performance over CL tasks.
@@
  • We propose a concise CL approach (TPPT) for CLIP based on incremental prompt tuning, without relying on complex regularization, external priors, or additional architectural overhead. TPPT leverages CLIP’s multi-modal nature and the stability of textual representations to explore its underutilized potential in CL settings.
  • We propose two variants of TPPT. TPPT-V learns visual prompts guided by fixed textual prototypes, which serve as stable anchors in the embedding space—capturing both class semantics and past knowledge beyond mere classification. To further bridge the vision-text gap in CL, TPPT-VT jointly learns visual and textual prompts while applying a diversity regularization to prevent embedding collapse and mitigate forgetting.
@@ 2 Related Work
Continual learning. Experience replay (ER) based methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022) involve storing a subset of the original training data and replaying it during the training of new classes to remind the model of old knowledge. Knowledge distillation-based methods (Li & Hoiem, 2017; Rebuffi et al., 2017; Douillard et al., 2020) use a teacher-student approach, where the original teacher model helps in training the updated student model by transferring knowledge, thus preserving previously learned information. Parameter regularization methods (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) add constraints to the learning process to retain important parameters from changing significantly, preserving prior knowledge. Dynamic networks (Yan et al., 2021; Wang et al., 2022a; b; Zhou et al., 2022a; Wang et al., 2024) adjust their architecture dynamically during learning, allowing different parts of the network to specialize in each task, thus balancing the learning of new classes with the preservation of old ones.
@@
Prompt tuning. Prompt tuning, originally proposed in the Natural Language Processing (NLP) domain (Lester et al., 2021; Liu et al., 2022; 2021), involves adapting pre-trained foundational models for downstream tasks. This is achieved by introducing a small number of learnable embeddings at the input, known as prompt tokens. Prompt tuning technique have demonstrated their effectiveness at adapting to downstream vision tasks as well (Zhou et al., 2022c; b; Yao et al., 2023; Jia et al., 2022; Khattak et al., 2023a; Rasheed et al., 2023; Khattak et al., 2023b), by prepending learnable prompt tokens to either the text encoder (Zhou et al., 2022c; b; Yao et al., 2023) or the image encoder (Jia et al., 2022), or learning multi-modal prompts (Rasheed et al., 2023; Khattak et al., 2023a; b) that are integrated into both text and image encoders. More recently, multi-modal prompt tuning (Rasheed et al., 2023; Khattak et al., 2023a; b) has emerged, which involves learning deep multi-modal prompts that jointly adapt the vision and language branches of the CLIP model for enhanced performance.
@@
CL based on prompt tuning. In light of the effectiveness of prompt tuning in adapting pre-trained models to downstream tasks, various prompt tuning-based methods (Wang et al., 2022e; d; c; 2023; Smith et al., 2023) have been developed for continual learning. Generally, these methods employ a set of prompts to preserve previously acquired knowledge, with each component designed to encode a segment of incoming knowledge. Specifically, L2P (Wang et al., 2022e) utilizes a fixed number of prompt sets and selects the top-K nearest prompts through query-key matching. DualPrompt (Wang et al., 2022d) innovates with both global prompts and an incremental set of task-specific prompts, also employing query-key matching for selection. S-Prompts (Wang et al., 2022c) adopts a similar set of prompt designs, but they are specifically tailored for domain-incremental learning. CODA-P (Smith et al., 2023) learns an incremental set of visual prompts. It employs a weighted aggregation of visual prompts and learns the prompts relying on orthogonality regularization.
@@ 3.1 Preliminaries
Class-incremental continual learning. Class-Incremental Learning (CIL) is a challenging scenario in the domain of CL (De Lange et al., 2021; Gomes et al., 2017; Mai et al., 2022), where a learning system is required to learn new classes over time without forgetting previously acquired knowledge (Lopez-Paz & Ranzato, 2017; Rebuffi et al., 2017). Formally, the learning process in CIL can be divided into a sequence of tasks ${\mathcal{D}}=\{{\mathcal{D}}^{1},{\mathcal{D}}^{2},\cdots,{\mathcal{D}}^{T}\}$, where each task ${\mathcal{D}}^{t}=\{({\mathbf{x}}_{i}^{t},y_{i}^{t})\}_{i=1}^{n_{t}}$ contains tuples of the input sample ${\mathbf{x}}_{i}^{t}\in{\mathcal{X}}$ and its corresponding label $y_{i}^{t}\in{\mathcal{Y}}$. Following conventional CIL settings (Rebuffi et al., 2017; Hou et al., 2019; Wu et al., 2019), a small number of exemplars from the seen classes are selected as the exemplar set $\mathcal{E}$}, and are used jointly with training data ${\mathcal{D}}_{t}$ of each task. CIL aims to develop a model that predicts the label $y$ given an unseen test sample $\mathbf{x}$} from arbitrary $T$ tasks.
CLIP model. In the CLIP model (Radford et al., 2021), the image and text encoders are represented as $f_{\theta}$} and $g_{\psi}$}, respectively. For an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, $C$ represent the height, width, and channel of the image, it undergoes segmentation into patches and subsequent processing through several Transformer blocks. This process yields a latent visual feature representation ${\mathbf{z}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. $D$ represents the dimension of the embedding space. The corresponding class label y{y} is integrated into hand-crafted prompts like “a photo of a [CLS]” to obtain text input $\mathbf{t}$}, where [CLS] denotes the class name associated with class label $y$. This prompt is then encoded by the text encoder $g_{\psi}$ to produce the text embedding ${\mathbf{w}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. The probability of a given image $\mathbf{x}$} being classified into class $y_{i}\in\{1,2,\cdots,C\}$ is calculated as follows:
@@
Prompt tuning for CLIP. Prompt tuning has emerged as a highly effective and efficient technique for adapting pre-trained foundational models to various downstream tasks. A range of methods (Zhou et al., 2022c; Yao et al., 2023; Rasheed et al., 2023; Khattak et al., 2023a; b) have been developed for adapting VLM. Specifically, a visual prompt ${\mathbf{P}}_{v,1}\in\mathbb{R}^{L_{v}\times D}$ and a textual prompt ${\mathbf{P}}_{t,1}\in\mathbb{R}^{L_{t}\times D}$ are prepended to the the visual and textual input tokens, steering the intermediate hidden states of the first layer of the model, and subsequently influence the final prediction. Here, $L_{v}$ and $L_{t}$ denote the lengths of the visual and textual prompts. The prompted visual and textual features are then denoted as $\tilde{{\mathbf{z}}}=\tilde{f_{\theta}}({\mathbf{x}})$ and $\tilde{{\mathbf{w}}}=\tilde{g_{\psi}}({\mathbf{t}})$, respectively. Here, $\tilde{f_{\theta}}$ and $\tilde{g_{\psi}}$ represent the encoders augmented with prompt tokens for image and text, respectively.
Figure 3: The overall framework of our 2 proposed methods. (1) The learned visual representations are guided by static textual prototypes (TPPT-V). We alleviate the forgetting issue by guiding visual representations with consistent textual prototypes, preventing drift of representations in the embedding space. (2) To improve upon the static textual prototypes, we propose to learn textual prompts for prototypes (TPPT-VT), and regulate the learning process by encouraging diversity. (3) Benefiting from the textual prototype anchors, our proposed methods remain simple yet effective, unlike previous methods that propose delicate, complex designs.

### v3

@@ Abstract
Continual learning (CL) enables deep neural networks to acquire new knowledge over time while mitigating catastrophic forgetting of previously learned information. The powerful generalization ability of pre-trained models (PTMs), such as the Contrastive Language-Image Pre-training (CLIP) model, has inspired a range of CL methods targeting new and specialized tasks, further bridging the gap between PTMs and continual adaptation. Leveraging its multi-modal visual and textual representations, CLIP offers a natural paradigm for CL, where new tasks can be accommodated by incrementally learning lightweight parameters, particularly prompts. However, existing prompt-based CL methods for PTMs often rely on complex designs built upon specific assumptions, such as intricate regularization schemes for prompt pools, specialized routing mechanisms, or multi-stage incrementation processes. While these approaches improve performance, they frequently introduce additional—and possibly unnecessary—complexity, underutilizing CLIP’s intrinsic capabilities. In this paper, we propose a concise CL approach for CLIP based on incremental prompt tuning that fully exploits its multi-modal structure and the stability of textual representations. Our method, Textual Prototype-guided Prompt Tuning (TPPT), introduces textual prototypes not merely as static classifiers, as in existing methods, but as stable anchors to guide the learning of visual prompts, thereby shaping the embedding space (i.e., TPPT-V). We show that our bidirectional supervision strategy enables more effective learning of new knowledge while reducing forgetting. To further close the vision-language gap during CL, we activate the language branch and extend our approach to jointly optimize both visual and textual prompts (i.e., TPPT-VT). We also introduce a relational diversity regularization on the textual anchors to prevent embedding space collapse and mitigate correlated forgetting. Extensive experiments and analyses demonstrate the effectiveness of our proposed approach, highlighting the benefits of leveraging CLIP’s intrinsic guidance for continual adaptation. Code is available at https://github.com/jeff024/tppt.
@@ 1 Introduction
Continual learning (CL) focuses on learning from a sequence of data or tasks while accumulating knowledge over time and avoiding catastrophic forgetting of previously learned information. It aims to do so with minimal storage and computational overhead by eliminating the need to retain all previously encountered data (de2021continual; gomes2017survey; mai2022online). Our primary focus in this study is class-incremental learning (CIL), a representative CL setting where classification tasks arrive incrementally, each consisting of distinct, non-overlapping classes (zhou2023deep; icarl). To mitigate catastrophic forgetting, various CL approaches have been explored, including experience replay (ER) methods (luo2023class; aljundi2019gradient; chaudhry2018riemannian; liu2020mnemonics; chaudhry2018efficient; yan2022learning) and regularization-based techniques (kirkpatrick2017overcoming; aljundi2018memory; zenke2017continual; aljundi2019task; npcl), both of which have shown notable success.
@@
Recent developments in pre-trained models (PTMs) based on the Transformer architecture (vaswani2017attention) have demonstrated strong generalization capabilities and achieved state-of-the-art performance across various applications (dosovitskiy2020image; devlin2018bert; wortsman2022robust). Among these, multi-modal pre-trained models, such as Contrastive Language-Image Pre-training (CLIP) (clip), jointly learn vision-language embedding spaces, effectively aligning visual and textual representations and enabling robust zero-shot generalization. To better adapt PTMs to downstream tasks, parameter-efficient fine-tuning methods such as prompt tuning (coop; kgcoop; maple; promptsrc) and adapters (gao2023clip; zhang2021tip; sung2022vl) have been developed.
@@
Figure 1: Conceptual illustrations of: (a) standard Cross-Entropy (CE), (b) our proposed TPPT-V, (c) a naïve multi-modal extension of TPPT-V, and (d) our proposed TPPT-VT. (a) Prior methods (coop; vpt; l2p; dualprompt; codaprompt; attriclip; proof) use CE loss to adapt PTMs, but suffer from representation drift (gama2014survey; lu2018learning), leading to forgetting. (b) TPPT-V introduces a textual prototypical contrastive loss to anchor visual features and mitigate drift. (c) A naïve extension that also tunes textual prompts may improve textual prototype quality but risks collapse to trivial solutions (minderer2205simple; kim2022transferring; maple; liang2022mind). (d) TPPT-VT addresses this by regularizing multi-modal prompt learning with diversity constraints on textual prototypes.
In CL settings, PTMs such as Vision Transformer (ViT) (dosovitskiy2020image) and CLIP (clip) often serve as stable, “frozen” backbones for continually arriving downstream tasks (e.g., in CIL). CL methods built on PTMs incrementally learn a fixed or dynamically expanding set of trainable parameters, such as soft prompts (l2p; dualprompt; lester2021power; liu2021p; codaprompt) or adapters (jha2024clap4clip; zhou2023revisiting; mcdonnell2024ranpac; sema; lu2025adaptiverankreducedforgetting; boosting), to incorporate new tasks encountered in the data stream. CLIP, with its jointly trained vision and language representations, provides a natural foundation for CL (jha2024clap4clip; attriclip), where textual embeddings can serve as pre-trained universal classifier heads for visual tasks. Due to their lightweight and non-intrusive nature, prompt-based methods have gained popularity in CL, particularly for CIL.
However, existing prompt-based continual learning methods with pre-trained models (including CLIP and ViT) often rely on task-specific prompt expansion (codaprompt; dualprompt), complex regularization using external priors like orthogonality (codaprompt; l2p; attriclip), or intricate prompt routing schemes (codaprompt; dualprompt; sprompts). While these methods can boost performance, they often introduce additional and potentially unnecessary complexity, potentially underutilizing the intrinsic strengths of CLIP’s multi-modal representations and the stability of its textual embeddings.
@@
It accommodates new tasks by incrementally adding visual prompts (vpt) to the inputs of multiple multi-head self-attention (MSA) layers within the CLIP encoders. While prompt tuning has been widely used for CL with PTMs such as CLIP (jha2024clap4clip) and ViT (codaprompt; dualprompt), prior works often introduce complexity through orthogonal regularization (codaprompt), probabilistic modeling (jha2024clap4clip), or handcrafted prompt selection (codaprompt; attriclip).
@@
We first present TPPT-V, which applies incremental prompt tuning only to the vision encoder of CLIP, guided by fixed textual embeddings of class labels. Prior works (attriclip; continualclip; ding2022don; zheng2023preventing) typically treat these textual embeddings as static classifiers and use cross entropy (CE) loss to align vision features with corresponding textual embeddings. While this is effective when all classes are jointly trained, relying solely on CE leads to representation drift and interference (gama2014survey; lu2018learning) in CL due to partial class visibility and the model’s unawareness of future tasks—resulting in forgetting (nguyen2019toward; mccloskey1989catastrophic), as illustrated in Fig. 1(a) and Fig. LABEL:fig:intro_drift.
As conceptualized in Fig. 1(a), this issue partly arises because CE loss focuses only on assigning each sample in isolation to a class label, ignoring how the embedding space evolves over time. We show that treating the fixed textual embeddings as prototype anchors (pcl; snell2017prototypical; ourpalm) and reintroducing contrastive learning supervision (clip; khosla2020supervised; grill2020bootstrap) alleviates this problem. Although the textual prototypes remain fixed, the contrastive loss (specifically, the asymmetric CE formulation) serves as an additional regularizer, guiding visual embeddings away from unrelated prototypes and stabilizing the embedding space (Fig. 1(b)). In this way, CLIP’s stable textual prototypes serve as anchors that help preserve prior knowledge by regularizing representation drift in the embedding space during CL.
(a) Analysis of (a) representation drift (lower is better) and (b) feature embedding diversity (higher is better to prevent collapse). In (a), we measure how far Task 1 embeddings deviate after learning all tasks. Training with CE loss alone leads to significant drift (Fig. 1(a)), whereas our proposed $\mathcal{L}_{\text{TPCL}}$ mitigates this (Fig. 1(b)). In (b), we assess embedding diversity via average pairwise distances over incremental stages. The red dotted line represents the diversity of the pre-trained CLIP model, denoted as ZS-CLIP. Naïve multi-modal training reduces diversity, especially on CUB, resulting in lower scores and potential collapse (Fig. 1(c)). Our diversity-regularized approach alleviates this, as shown in Fig. 1(d).
To further enhance vision-language alignment in CL on downstream tasks, we go beyond vision-only tuning and enable learning of both visual and textual prompts. However, naïve joint prompt tuning can cause embedding space collapse (minderer2205simple; kim2022transferring; liang2022mind; maple) (as shown in Fig. 1(c) and Fig. LABEL:fig:intro_collapse), especially in CL, where the model can easily converge to trivial solutions (minderer2205simple; kim2022transferring; liang2022mind; maple). To prevent this, we propose TPPT-VT (illustrated in Fig. 1(d)), which introduces relational diversity regularization on textual prompts, maintaining diversity among textual prototypes and mitigating collapse. By preserving the original contextual structure of textual representations, TPPT-VT improves performance over CL tasks.
@@
  • We propose a concise CL approach (TPPT) for CLIP based on incremental prompt tuning, without relying on complex regularization, external priors, or additional architectural overhead. TPPT leverages CLIP’s multi-modal nature and the stability of textual representations to explore its underutilized potential in CL.
  • We propose two variants of TPPT. TPPT-V learns visual prompts guided by fixed textual prototypes, which serve as stable anchors in the embedding space—capturing both class semantics and past knowledge beyond mere classification. To further bridge the vision-text gap in CL, TPPT-VT jointly learns visual and textual prompts while applying a diversity regularization to prevent embedding collapse and forgetting.
@@ 2 Related Work
Continual learning. Experience replay (ER) based methods (luo2023class; aljundi2019gradient; chaudhry2018riemannian; liu2020mnemonics; chaudhry2018efficient; yan2022learning; tong2025coreset) involve storing a subset of the original training data and replaying it during the training of new classes to remind the model of old knowledge. Knowledge distillation-based methods (lwf; icarl; douillard2020podnet) use a teacher-student approach, where the original teacher model helps in training the updated student model by transferring knowledge, thus preserving previously learned information. Parameter regularization methods (kirkpatrick2017overcoming; aljundi2018memory; zenke2017continual; aljundi2019task; npcl) add constraints to the learning process to retain important parameters from changing significantly, preserving prior knowledge. Dynamic networks (der; wang2022beef; wang2022foster; zhou2022model; sema) adjust their architecture dynamically during learning, allowing different parts of the network to specialize in each task, thus balancing the learning of new classes with the preservation of old ones.
@@
Prompt tuning. Prompt tuning, originally proposed in the Natural Language Processing (NLP) domain (lester2021power; liu2022p; liu2021p), involves adapting pre-trained foundational models for downstream tasks. This is achieved by introducing a small number of learnable embeddings at the input, known as prompt tokens. Prompt tuning technique have demonstrated their effectiveness at adapting to downstream vision tasks as well (coop; cocoop; kgcoop; vpt; maple; ivlp; promptsrc), by prepending learnable prompt tokens to either the text encoder (coop; cocoop; kgcoop) or the image encoder (vpt), or learning multi-modal prompts (ivlp; maple; promptsrc) that are integrated into both text and image encoders. More recently, multi-modal prompt tuning (ivlp; maple; promptsrc) has emerged, which involves learning deep multi-modal prompts that jointly adapt the vision and language branches of the CLIP model for enhanced performance.
@@
CL based on prompt tuning. In light of the effectiveness of prompt tuning in adapting pre-trained models to downstream tasks, various prompt tuning-based methods (l2p; dualprompt; sprompts; attriclip; codaprompt) have been developed for continual learning. Generally, these methods employ a set of prompts to preserve previously acquired knowledge, with each component designed to encode a segment of incoming knowledge. Specifically, L2P (l2p) utilizes a fixed number of prompt sets and selects the top-K nearest prompts through query-key matching. DualPrompt (dualprompt) innovates with both global prompts and an incremental set of task-specific prompts, also employing query-key matching for selection. S-Prompts (sprompts) adopts a similar set of prompt designs, but they are specifically tailored for domain-incremental learning. CODA-P (codaprompt) learns an incremental set of visual prompts. It employs a weighted aggregation of visual prompts and learns the prompts relying on orthogonality regularization.
@@ 3.1 Preliminaries
Class-incremental continual learning. Class-Incremental Learning (CIL) is a challenging scenario in the domain of CL (de2021continual; gomes2017survey; mai2022online), where a learning system is required to learn new classes over time without forgetting previously acquired knowledge (lopez2017gradient; icarl). Formally, the learning process in CIL can be divided into a sequence of tasks ${\mathcal{D}}=\{{\mathcal{D}}^{1},{\mathcal{D}}^{2},\cdots,{\mathcal{D}}^{T}\}$, where each task ${\mathcal{D}}^{t}=\{({\mathbf{x}}_{i}^{t},y_{i}^{t})\}_{i=1}^{n_{t}}$ contains tuples of the input sample ${\mathbf{x}}_{i}^{t}\in{\mathcal{X}}$ and its corresponding label $y_{i}^{t}\in{\mathcal{Y}}$. Following conventional CIL settings (icarl; hou2019learning; wu2019large), a small number of exemplars from the seen classes are selected as the exemplar set $\mathcal{E}$}, and are used jointly with training data ${\mathcal{D}}_{t}$ of each task. CIL aims to develop a model that predicts the label $y$ given an unseen test sample $\mathbf{x}$} from arbitrary $T$ tasks.
CLIP model. In the CLIP model (clip), the image and text encoders are represented as $f_{\theta}$} and $g_{\psi}$}, respectively. For an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, $C$ represent the height, width, and channel of the image, it undergoes segmentation into patches and subsequent processing through several Transformer blocks. This process yields a latent visual feature representation ${\mathbf{z}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. $D$ represents the dimension of the embedding space. The corresponding class label y{y} is integrated into hand-crafted prompts like “a photo of a [CLS]” to obtain text input $\mathbf{t}$}, where [CLS] denotes the class name associated with class label $y$. This prompt is then encoded by the text encoder $g_{\psi}$ to produce the text embedding ${\mathbf{w}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. The probability of a given image $\mathbf{x}$} being classified into class $y_{i}\in\{1,2,\cdots,C\}$ is calculated as follows:
@@
Prompt tuning for CLIP. Prompt tuning has emerged as a highly effective and efficient technique for adapting pre-trained foundational models to various downstream tasks. A range of methods (coop; kgcoop; ivlp; maple; promptsrc) have been developed for adapting VLM. Specifically, a visual prompt ${\mathbf{P}}_{v,1}\in\mathbb{R}^{L_{v}\times D}$ and a textual prompt ${\mathbf{P}}_{t,1}\in\mathbb{R}^{L_{t}\times D}$ are prepended to the the visual and textual input tokens, steering the intermediate hidden states of the first layer of the model, and subsequently influence the final prediction. Here, $L_{v}$ and $L_{t}$ denote the lengths of the visual and textual prompts. The prompted visual and textual features are then denoted as $\tilde{{\mathbf{z}}}=\tilde{f_{\theta}}({\mathbf{x}})$ and $\tilde{{\mathbf{w}}}=\tilde{g_{\psi}}({\mathbf{t}})$, respectively. Here, $\tilde{f_{\theta}}$ and $\tilde{g_{\psi}}$ represent the encoders augmented with prompt tokens for image and text, respectively.
Figure 3: The overall framework of our 2 proposed methods. (1) The learned visual representations are guided by static textual prototypes (TPPT-V). We alleviate the forgetting issue by guiding visual representations with consistent textual prototypes, preventing drift of representations in the embedding space. (2) To improve upon the static textual prototypes, we propose to learn textual prompts for prototypes (TPPT-VT), and regulate the learning process by encouraging diversity. (3) Benefiting from the textual prototype anchors, our proposed methods remain simple yet effective, unlike previous methods that use delicate, complex designs.